# Strong Normalization for the Safe Fragment of a Minimal Rewrite System: A Triple-Lexicographic Proof and a Conjecture on the Unprovability of Full Termination for Any Relational Operator-Only TRS

### v1

@@ Abstract
We present a minimal operator-only term rewriting system with seven constructors and eight reduction rules. Our main contribution is a mechanically-verified proof of strong normalization for a guarded fragment using a novel triple-lexicographic measure combining a phase bit, multiset ordering (Dershowitz-Manna), and ordinal ranking. From strong normalization, we derive a certified normalizer with proven totality and soundness. Assuming local confluence (verified through critical pair analysis), Newman’s Lemma yields confluence and therefore unique normal forms for the safe fragment. We establish impossibility results showing that simpler measures, such as additive counters, polynomial interpretations, and single-bit flags, provably fail for rules with term duplication. The work demonstrates fundamental limitations in termination proving for self-referential systems. We state a conjecture: no relational operator-only TRS can have its full-system termination proved by internally definable methods. Here “relational” is equivalent to “capable of ordered computation”: systems with a recursor enabling iteration over successors, comparison, or sequential counting. Such recursors necessarily redistribute step arguments across recursive calls, defeating all additive termination measures. This structural limitation applies to any TRS expressive enough to encode ordered data. All theorems have been formally verified in a proof assistant. The formal development is available to program committee members and referees upon request for purposes of peer review.
@@ 1 Introduction
We develop a minimal operator-only rewrite calculus (KO7): the object language contains only constructors and operators with rewrite rules; there are no binders, types, external axioms, or semantic predicates. The rules are the semantics. Our goals are: (i) a clean, duplication-robust proof of strong normalization (SN); (ii) a certified normalizer that always returns a normal form; (iii) (optionally) unique normal forms via Newman’s Lemma under a local-confluence assumption; and (iv) a conjecture that no operator-only TRS capable of ordered computation can have its full-system termination proved by internally definable methods.
Scope. All formal results are established for a guarded safe subrelation. We do not claim a single global measure for the full unguarded relation. Moreover, we exhibit a precise negative: at the root peak $\mathrm{eqW}\,a\,a$ with $\kappa^{M}(a)=0$, local join fails in the full relation. Therefore the full relation is not locally confluent at that peak, hence not confluent. The designed remedy is to work in the safe fragment with guarded/context joiners (§6).
A second conceptual goal is to situate KO7 against results about fixed-target reachability in terminating TRSs. If we define an internal provability predicate by “$t$ reduces to $\top$”, then under SN the set $\{t\mid t\Rightarrow^{\ast}\top\}$ is decidable by normalization/backtracking. With confluence, decision reduces to normal-form equality. This explains why single-level Gödel encodings cannot coexist with globally terminating proof search; the right move is stratification.
@@
(1) A duplication-robust SN proof for KO7 using a triple-lex measure with a multiset (DM) component and an MPO-style head precedence; (2) a total, proved-correct normalizer; (3) a guarded Newman module for the safe relation that yields confluence (hence unique normal forms) from SN + local confluence; (4) a decidability result for reachability under SN; (5) a catalog of impossibility results for additive and polynomial measures under duplication; (6) a formal verification of all results; (7) a conjecture on undecidable termination for operator-only systems within a fixed base theory.
@@
  • SN (SafeStep) via triple lex: Formally proven using a lexicographic measure combining a $\delta$-phase bit, a Dershowitz–Manna multiset rank, and an ordinal.
@@
  • Full Step caveat: We exhibit a specific peak ($\mathrm{eqW}\,a\,a$ with $\kappa^{M}(a)=0$) where local join fails, justifying the restriction to the safe subrelation.
@@ 2 Background: TRSs, SN, reachability, and Newman
We assume standard abstract reduction and term rewriting notions [BaaderNipkow98, Terese03]. A term is in normal form if no rule applies. A TRS is strongly normalizing (SN) if there are no infinite reductions. A relation is confluent if for any $t\Rightarrow^{\ast}u$ and $t\Rightarrow^{\ast}v$ there exists $w$ with $u\Rightarrow^{\ast}w$ and $v\Rightarrow^{\ast}w$. Local confluence requires this only for single-step forks. Newman’s Lemma asserts SN + local confluence $\Rightarrow$ confluence [Newman42], yielding unique normal forms.
Fixed-target (“small-term”) reachability in terminating TRSs has well-charted complexity: NP-complete for length-reducing systems (dropping to P under confluence), NExpTime/N2ExpTime for (linear) polynomial interpretations, and PSPACE for KBO-terminating TRSs [BaaderGiesl24]. Modularity holds in certain linear, non-collapsing combinations [CaronCoquide94], while even flat non-linear systems exhibit undecidable reachability and confluence [Mitsuhashi06]. Termination under duplication typically requires orders beyond plain sizes; a naive size can increase under duplicating rules. The cure is a multiset extension of a base order [DershowitzManna79] or (recursive) path orders [Dershowitz87].
@@ 3 The KO7 calculus
KO7 is a finite TRS over a small signature (7 constructors) with 8 rules (including an explicit case split on eqW). For concreteness we list the rule shapes below.
@@
R1 merge 2 $\mathrm{merge}\;\mathrm{void}\;t\to t$ No
R2 merge 2 $\mathrm{merge}\;t\;\mathrm{void}\to t$ No
R3 merge 2 $\mathrm{merge}\;t\;t\to t$ No
@@
R6 integrate 1 $\mathrm{integrate}\,(\mathrm{delta}\,t)\to\mathrm{void}$ No
R7 eqW 2 $\mathrm{eqW}\;a\;a\to\mathrm{void}$ No
R8 eqW 2 $\mathrm{eqW}\;a\;b\to\mathrm{integrate}\,(\mathrm{merge}\,a\,b)$ (if $a\neq b$) No
Table 1: KO7’s 8 rules. R7/R8 provide a complete case split on equality. R3 is collapsing. R5 redistributes $s$ to two RHS positions.
@@
Note. We prove SN and Newman-based confluence for the safe fragment.
@@ 4 Strong normalization
We define a triple-lexicographic measure
@@
ordered by the lex product of: (i) a phase bit dropping on the successor recursion; (ii) a multiset of ranks $\kappa^{M}$ (Dershowitz–Manna) with an explicit precedence/status orienting redex >> pieces; and (iii) an ordinal payload $\mu_{\mathrm{ord}}$ for non-duplicating ties. Ordinal hazards are stated explicitly (right-addition is not strictly monotone; absorption $\alpha+\beta=\beta$ requires $\omega\leq\beta$). In duplicating branches (if extended to broader systems) we would use a compact MPO-style head precedence.
@@
For the SafeStep relation we use a single unified triple lex order $(\delta,\kappa^{M},\mu_{\mathrm{ord}})$. For the full relation we only provide a disjunctive decrease certificate: each step is covered by either a KO7 lex drop or an MPO triple drop. We do not claim a single global well-founded measure for all full steps.
@@
We record minimal counterpatterns that motivate the multiset/path components:
@@ 5 A certified normalizer
By well-founded recursion on $\mu^{3}$ we define a normalization function. We prove the following properties (separated to avoid overfull lines):
@@
These are exposed in the formalization. We make no efficiency claim: worst-case normalization cost follows the termination witness in play (cf. the small-term reachability bounds in [BaaderGiesl24]). In confluent systems, decision reduces to one normalization and an equality check.
@@ 6 Local confluence and Newman (guarded safe relation)
We discharge local confluence by joining the finite set of critical pairs (when present). Combining Cor. 1 with Newman’s Lemma [Newman42] yields:
@@
We work with a guarded safe subrelation SafeStep for which we prove:
@@
  • Local Confluence (SafeStep). We provide local-join lemmas per root shape and context wrappers. Newman then yields confluence and unique NFs for the safe fragment.
@@
The following table covers the safe root configurations with explicit local-join lemmas. This is exhaustive for SafeStep except the reflexive $\mathrm{eqW}\,a\,a$ peak under $\kappa^{M}(a)=0$, which we show is not locally joinable at the root. We discharge many eqW cases via guarded/context wrappers.
@@
Define the safe-phase predicate by $\delta\!\operatorname{-guard}(t)\iff\delta\!\operatorname{Flag}(t)=0$. Here $\delta\!\operatorname{Flag}:\mathrm{Term}\to\mathbb{N}$ is a structurally recursive function defined on terms in the artifact (tracking split parity), so the predicate is decidable. Facts: $\delta\!\operatorname{Flag}(\mathrm{eqW}\,a\,b)=0$; merge-void rules require $\delta\!\operatorname{Flag}(t)=0$.
@@ 7 Impossibility results
We formally establish the failure of several simpler measure strategies, which necessitate the use of DM multiset orders or MPO. These negative results are verified in the formal development.
@@
We use a simple contract for internally defined termination measures to structure these negative results:
@@
An internally definable measure for a type $\alpha$ consists of $(\beta,\;<\_{\beta},\;\mathsf{wf},\;m,\;\mathsf{ctxMono},\;\mathsf{piecesLt})$ where: $\beta$ is a base order carrier; $_{\beta}$ is a well-founded relation with witness $\mathsf{wf}$; $m:\alpha\to\beta$ is the measure; $\mathsf{ctxMono}$ expresses context compatibility; and $\mathsf{piecesLt}$ asserts that in each rule instance, every RHS piece is strictly smaller than the removed LHS redex w.r.t. $_{\beta}$.
@@ 8 Decidability of Reachability
This connects our work to classical results: if we could encode undecidable properties through reduction to constants, we would violate known theoretical limits. This motivates stratified approaches in proof assistants, where object-level and meta-level reasoning are carefully separated.
@@
We work with a finite first-order TRS over finite terms. The decision method is: compute a normal form (by SN) and check whether it is $\top$; with local confluence, this reduces to one normalization and an equality test.
@@
Small-term reachability in terminating TRSs ranges from NP (length-reducing) to NExpTime/N2ExpTime (polynomial interpretations) and PSPACE (KBO), with confluence lowering the length-reducing class to P [BaaderGiesl24]. This situates our “normalize and compare” decision procedure for KO7 within the established landscape.
Moreover, decidability can be modular for disjoint unions under left-linearity/non-collapsing assumptions [CaronCoquide94], while termination alone does not guarantee decidability: even flat non-linear TRSs have undecidable reachability/joinability/confluence [Mitsuhashi06].
@@ 10 Formalization structure
All claimed results are formally proven without reliance on unproven postulates in the current build. The formal development is available to reviewers under appropriate confidentiality agreements.
@@
KO7 Rules Table (7 constructors, 8 rules)
Rule Head Arity Shape Dup?
R1 merge 2 $\mathrm{merge}\;\mathrm{void}\;t\to t$ No
R2 merge 2 $\mathrm{merge}\;t\;\mathrm{void}\to t$ No
R3 merge 2 $\mathrm{merge}\;t\;t\to t$ No
[…]
R6 integrate 1 $\mathrm{integrate}\,(\mathrm{delta}\,t)\to\mathrm{void}$ No
R7 eqW 2 $\mathrm{eqW}\;a\;a\to\mathrm{void}$ No
R8 eqW 2 $\mathrm{eqW}\;a\;b\to\mathrm{integrate}\,(\mathrm{merge}\,a\,b)$ (if $a\neq b$) No
Table 2: KO7’s 8 rules. R7/R8 provide a complete case split on equality. R3 is collapsing (erases one copy). R5 redistributes $s$ to two positions on the RHS; see Remark below.
@@
We order $\mu^{3}(t)=(\delta\text{-flag}(t),\,\kappa^{M}(t),\,\mu_{\mathrm{ord}}(t))$ by lex:
@@
  • Multiset $\kappa^{M}$: a Dershowitz–Manna multiset extension over a base precedence/status (or MPO head precedence) orienting redexes strictly above pieces; covers duplicators.
@@
We emphasize that $\kappa^{M}$ aggregates via multiset union ($\cup$), not numeric addition. For duplicating rules, the multiset of piece-weights on the RHS is DM-smaller than the singleton multiset containing the LHS redex weight, yielding a strict drop in the $\kappa^{M}$ component. For non-duplicating rules, $\kappa^{M}$ ties by definitional equality and the ordinal $\mu_{\mathrm{ord}}$ resolves the branch. In particular, for unguarded instances of eqW-refl and merge-cancel we use a $\kappa$-branch: when $\kappa^{M}(a)\neq 0$ we obtain a left-lex drop via DM; when $\kappa^{M}(a)=0$ the $\kappa^{M}$ component ties by rfl and the strict decrease is witnessed in the $\mu_{\mathrm{ord}}$ coordinate (right branch of Prod.Lex).
@@
SN + local confluence implies confluence; in the artifact this is instantiated for the safe relation via an Acc-based star–star join. Scope note: all confluence statements and their Newman instantiations are for SafeStep only. Full Step is not locally joinable at root for eqW a a with $\kappa^{M}(a)=0$; accordingly we do not claim full-step confluence.
@@
merge void left $\mu$ (ordinal) Theorem 4.1 Termination
merge void right $\mu$ (ordinal) Theorem 4.2 Termination
@@
rec succ $\delta$ phase bit (1 $\to$ 0) Theorem 4.5 Termination
@@
integrate (delta t) $\mu$ (unique target) Theorem 4.8 Termination
toy duplication DM on size s >> pair Lemma 7.1 Impossibility
toy duplication MPO triple headRank(s) >> headRank(pair) Lemma 7.2 Impossibility
Table 3: Per-rule orientation summary.
Parenthetical aliases: for substitution convenience we expose simp forms alongside the main rec lemmas.
Local-join report (critical peaks and unique-target cases)
Shape Join lemma name Notes
merge void void localJoin_merge_void_void both branches to void
merge t t localJoin_merge_tt joins at $t$ (context-lifted)
integrate (delta t) localJoin_int_delta unique target void
@@
eqW a b, a $\neq$ b localJoin_eqW_ne only diff applies (unique target)
eqW a a, guard kappaM a $\neq$ 0 localJoin_eqW_refl_guard_ne refl blocked; only diff
merge a b (no void, a $\neq$ b) localJoin_merge_no_void_neq no root step; vacuous
integrate non-delta localJoin_integrate_non_delta no root step; vacuous
[…]
Table 4: Local-join witnesses for the safe relation (root).
Ctx-local join wrappers
When needed in context, the root local-join lemmas lift to context via wrappers:
Shape Ctx wrapper lemma Notes
merge a b (no void, a $\neq$ b) localJoin_ctx_merge_no_void_neq from root ‘merge’ lemma
eqW a b, a $\neq$ b localJoin_ctx_eqW_ne only diff applies
eqW a a, kappaM guard localJoin_ctx_eqW_refl_guard_ne refl blocked
[…]
eqW (delta n) (delta n) localJoin_ctx_eqW_refl_when_a_is_delta literal delta case
Table 5: Ctx-local join wrappers.
@@
This summarizes the exact orientation strategy realized in the proofs (no global RPO required):
@@
  • The rule rec_zero uses a DM multiset drop on $\kappa^{M}$, wrapped by drop_R_rec_zero.
  • The rule rec_succ drops the $\delta$ phase bit (1 $\to$ 0) via drop_R_rec_succ.
@@
  • Conceptual precedence (if one prefers an MPO reading) aligns with rec >> merge >> app >> integrate >> eqW >> delta >> void, but the formal proofs do not require a global precedence declaration.
Formal anchors:
  • Merge void left/right: $\mu$-first or DM tie plus $\mu$.
  • Merge cancel: DM multiset drop on $\kappa^{M}$.
  • Rec zero: drop_R_rec_zero with dm_drop_R_rec_zero for the inner DM piece.
  • Rec succ: drop_R_rec_succ via the $\delta$ phase bit (1 $\to$ 0).
  • EqW refl/diff: $\mu$-first MPO triple.
  • Integrate delta: unique target void.

### v2

@@ Abstract
This paper presents a minimal operator-only term rewriting system with seven constructors and eight reduction rules. The main contribution is a mechanically-verified proof of strong normalization for a guarded fragment using a novel triple-lexicographic measure combining a phase bit, multiset ordering (Dershowitz-Manna), and ordinal ranking. From strong normalization, a certified normalizer with proven totality and soundness is derived. Assuming local confluence (verified through critical pair analysis), Newman’s Lemma yields confluence and therefore unique normal forms for the safe fragment. Impossibility results showing that simpler measures, such as additive counters, polynomial interpretations, and single-bit flags, provably fail for rules with term duplication, are established. The work demonstrates fundamental limitations in termination proving for self-referential systems. A conjecture is stated: no relational operator-only TRS can have its full-system termination proved by internally definable methods. Here “relational” is equivalent to “capable of ordered computation”: systems with a recursor enabling iteration over successors, comparison, or sequential counting. Such recursors necessarily redistribute step arguments across recursive calls, defeating all additive termination measures. This structural limitation applies to any TRS expressive enough to encode ordered data. All theorems have been formally verified in a proof assistant. The Lean formalization is available at https://github.com/MosesRahnama/OperatorKO7.
@@ 1 Introduction
This paper presents a minimal operator-only rewrite calculus (KO7) where the object language contains only constructors and operators with rewrite rules. There are no binders, types, external axioms, or semantic predicates; the rules are the semantics. The primary goals are: (i) a clean, duplication-robust proof of strong normalization (SN); (ii) a certified normalizer that always returns a normal form; (iii) unique normal forms via Newman’s Lemma under a local-confluence assumption; and (iv) a conjecture regarding the unprovability of termination for relational operator-only systems using internal methods.
Scope. The formal results apply to a guarded safe subrelation. This restriction is necessary because the full relation is not locally confluent at the root peak $\mathrm{eqW}\,a\,a$ when $\kappa^{M}(a)=0$. The safe fragment employs guarded/context joiners to ensure unique normal forms (§6).
A second conceptual goal is to situate KO7 against results about fixed-target reachability in terminating TRSs. If an internal provability predicate is defined by “$t$ reduces to $\top$”, then under SN the set decidability of $\{t\mid t\to^{*}\top\}$ is reducible to normalization. With confluence, decision reduces to normal-form equality. This explains why single-level Gödel encodings cannot coexist with globally terminating proof search; stratification is required.
@@
Strong normalization is proven for a guarded fragment of KO7 using a novel triple-lexicographic measure that combines a phase bit, multiset ordering (Dershowitz-Manna), and ordinal ranking. This result supports a certified normalizer with proven totality and soundness. Furthermore, confluence for the safe fragment is established via Newman’s Lemma given verified local confluence. The work also includes formally verified impossibility results for additive and polynomial measures under duplication, demonstrating fundamental limitations in termination proving for self-referential systems.
@@
  • SN (SafeStep) via triple lex: Formally proven using a lexicographic measure combining a $\delta$-phase bit, a Dershowitz-Manna multiset rank, and an ordinal.
@@
  • Full Step caveat: A specific peak ($\mathrm{eqW}\,a\,a$ with $\kappa^{M}(a)=0$) where local join fails is exhibited, justifying the restriction to the safe subrelation.
@@ 2 Background: TRSs, SN, reachability, and Newman
Standard abstract reduction and term rewriting notions apply [2, 11]. A term is in normal form if no rule applies. A TRS is strongly normalizing (SN) if there are no infinite reductions. A relation is confluent if for any $t\Rightarrow^{\ast}u$ and $t\Rightarrow^{\ast}v$ there exists $w$ with $u\Rightarrow^{\ast}w$ and $v\Rightarrow^{\ast}w$. Local confluence requires this only for single-step forks. Newman’s Lemma asserts SN + local confluence $\Rightarrow$ confluence [10], yielding unique normal forms.
Fixed-target (“small-term”) reachability in terminating TRSs has well-charted complexity: NP-complete for length-reducing systems (dropping to P under confluence), NExpTime/N2ExpTime for (linear) polynomial interpretations, and PSPACE for KBO-terminating TRSs [1]. Modularity holds in certain linear, non-collapsing combinations [4], while even flat non-linear systems exhibit undecidable reachability and confluence [9]. Termination under duplication typically requires orders beyond plain sizes; a naive size can increase under duplicating rules. The cure is a multiset extension of a base order [5] or (recursive) path orders [6].
@@ 3 The KO7 calculus
KO7 is a finite TRS over a small signature (7 constructors) with 8 rules (including the equality-witness operator eqW). For concreteness we list the kernel rule shapes below.
@@
R_merge_void_left merge 2 $\mathrm{merge}\;\mathrm{void}\;t\to t$ No
R_merge_void_right merge 2 $\mathrm{merge}\;t\;\mathrm{void}\to t$ No
R_merge_cancel merge 2 $\mathrm{merge}\;t\;t\to t$ No
@@
R_int_delta integrate 1 $\mathrm{integrate}\,(\mathrm{delta}\,t)\to\mathrm{void}$ No
R_eq_refl eqW 2 $\mathrm{eqW}\;a\;a\to\mathrm{void}$ No
R_eq_diff eqW 2 $\mathrm{eqW}\;a\;b\to\mathrm{integrate}\,(\mathrm{merge}\,a\,b)$ No
Table 1: KO7’s 8 kernel rules. In the full kernel relation, R_eq_refl and R_eq_diff overlap at eqW a a. The certified safe fragment guards R_eq_diff by $a\neq b$. R_merge_cancel is collapsing. R_rec_succ redistributes $s$ to two RHS positions.
Kernel signature (7 constructors).
Terms are generated by the constructors:
$t::=\mathrm{void}\mid\delta\,t\mid\mathrm{integrate}\,t\mid\mathrm{merge}\,t\,t\mid\mathrm{app}\,t\,t\mid\mathrm{rec}\,\Delta\,t\,t\,t\mid\mathrm{eqW}\,t\,t.$
The full kernel relation Step is generated by the 8 rules in Table 1, while the certified artifact is proved for a guarded subrelation SafeStep.
@@
Note. Strong normalization and Newman-based confluence are proven for the safe fragment.
@@ 4 Strong normalization
The termination argument relies on a triple-lexicographic measure
@@
ordered by the lex product of: (i) a phase bit dropping on the successor recursion; (ii) a multiset of ranks $\kappa^{M}$ (Dershowitz-Manna) with an explicit precedence/status orienting redex >> pieces; and (iii) an ordinal payload $\mu_{\mathrm{ord}}$ for non-duplicating ties. Ordinal hazards are formally handled (right-addition is not strictly monotone; absorption $\alpha+\beta=\beta$ requires $\omega\leq\beta$). Duplicating branches use a compact MPO-style head precedence.
@@
For the SafeStep relation, a single unified triple lex order $(\delta,\kappa^{M},\mu_{\mathrm{ord}})$ suffices. Steps in the full relation are covered by a disjunctive decrease certificate: each step corresponds to either a KO7 lex drop or an MPO triple drop.
@@
Minimal counterpatterns that motivate the multiset/path components are recorded:
@@ 5 A certified normalizer
A normalization function is defined by well-founded recursion on $\mu^{3}$. Properties proven in the formalization include:
@@
No specific efficiency claim is made: worst-case normalization cost follows the termination witness (cf. the small-term reachability bounds in [1]). Under confluence, decision reduces to one normalization and an equality check.
@@ 6 Local confluence and Newman (guarded safe relation)
Local confluence is discharged by joining the finite set of critical pairs. Combining Cor. 1 with Newman’s Lemma [10] yields:
@@
The formal results concern a guarded safe subrelation SafeStep for which we prove:
@@
  • Local Confluence (SafeStep). Local-join lemmas per root shape and context wrappers are provided. Newman then yields confluence and unique NFs for the safe fragment.
@@
The following table covers the safe root configurations with explicit local-join lemmas. This is exhaustive for SafeStep except the reflexive $\mathrm{eqW}\,a\,a$ peak under $\kappa^{M}(a)=0$, which is shown to be not locally joinable at the root. We discharge many eqW cases via guarded/context wrappers.
@@
Define the safe-phase predicate by $\delta\text{-guard}(t)\iff\mathrm{deltaFlag}(t)=0$. Here $\mathrm{deltaFlag}:\mathrm{Term}\to\mathbb{N}$ is a structurally recursive function defined on terms in the artifact (tracking split parity), so the predicate is decidable. Facts: $\mathrm{deltaFlag}(\mathrm{eqW}\,a\,b)=0$; merge-void rules require $\mathrm{deltaFlag}(t)=0$.
@@ 7 Impossibility results
The formal development establishes the failure of several simpler measure strategies, necessitating the use of DM multiset orders or MPO:
@@
A simple contract for internally defined termination measures structures these negative results:
@@
An internally definable measure for a type $\alpha$ consists of $(\beta,\;<\_{\beta},\;\mathsf{wf},\;m,\;\mathsf{ctxMono},\;\mathsf{piecesLt})$ where: $\beta$ is a base order carrier; $_{\beta}$ is a well-founded relation with explicit witness $\mathsf{wf}$; $m:\alpha\to\beta$ is the measure; $\mathsf{ctxMono}$ expresses context compatibility; and $\mathsf{piecesLt}$ asserts that in each rule instance, every RHS piece is strictly smaller than the removed LHS redex w.r.t. $_{\beta}$.
@@ 8 Decidability of Reachability
This connects the work to classical results: encoding undecidable properties through reduction to constants violates known theoretical limits. This motivates stratified approaches in proof assistants where object-level and meta-level reasoning are kept separate.
@@
The model assumes a finite first-order TRS over finite terms. The decision procedure computes a normal form (by SN) and checks whether it is $\top$; with local confluence, this reduces to one normalization and an equality test.
@@
Small-term reachability in terminating TRSs ranges from NP (length-reducing) to NExpTime/N2ExpTime (polynomial interpretations) and PSPACE (KBO), with confluence lowering the length-reducing class to P [1]. This situates our “normalize and compare” decision procedure for KO7 within the established landscape.
Moreover, decidability can be modular for disjoint unions under left-linearity/non-collapsing assumptions [4], while termination alone does not guarantee decidability: even flat non-linear TRSs have undecidable reachability/joinability/confluence [9].
@@ 10 Formalization structure
All claimed results are formally proven without reliance on unproven postulates in the current build. The formal development is available at https://github.com/MosesRahnama/OperatorKO7.
Acknowledgments
AI has been a great help with this project.
@@ Acknowledgments
Lean kernel location
The authoritative mechanized kernel is the Lean file OperatorKO7/Kernel.lean in the repository https://github.com/MosesRahnama/OperatorKO7. The safe fragment and proofs are under OperatorKO7/Meta/.
@@
The measure $\mu^{3}(t)=(\delta\text{-flag}(t),\,\kappa^{M}(t),\,\mu_{\mathrm{ord}}(t))$ is ordered by lex:
@@
  • Multiset $\kappa^{M}$: a Dershowitz-Manna multiset extension over a base precedence/status (or MPO head precedence) orienting redexes strictly above pieces; covers duplicators.
@@
It is emphasized that $\kappa^{M}$ aggregates via multiset union ($\cup$), not numeric addition. For duplicating rules, the multiset of piece-weights on the RHS is DM-smaller than the singleton multiset containing the LHS redex weight, yielding a strict drop in the $\kappa^{M}$ component. For non-duplicating rules, $\kappa^{M}$ ties by definitional equality and the ordinal $\mu_{\mathrm{ord}}$ resolves the branch. In particular, for unguarded instances of eqW-refl and merge-cancel a $\kappa$-branch strategy is employed: when $\kappa^{M}(a)\neq 0$ a left-lex drop is obtained via DM; when $\kappa^{M}(a)=0$ the $\kappa^{M}$ component ties by rfl and the strict decrease is witnessed in the $\mu_{\mathrm{ord}}$ coordinate (right branch of Prod.Lex).
@@
SN + local confluence implies confluence; in the artifact this is instantiated for the safe relation via an Acc-based star-star join. Scope note: all confluence statements and their Newman instantiations are for SafeStep only. Full Step is not locally joinable at root for eqW a a with $\kappa^{M}(a)=0$; accordingly full-step confluence is not claimed.
@@
merge void left $\mu$ (ordinal) (n/a) Theorem 4.1 Termination
merge void right $\mu$ (ordinal) (n/a) Theorem 4.2 Termination
@@
rec succ $\delta$ phase bit (1 $\to$ 0) (n/a) Theorem 4.5 Termination
@@
integrate (delta t) $\mu$ (unique target) (n/a) Theorem 4.8 Termination
Table 2: Termination Proof Strategy: Per-rule Orientation Summary.
Parenthetical aliases: for substitution convenience simp forms are exposed alongside the main rec lemmas.
Local-join report (Confluence)
Type Shape Join lemma name Notes
Root merge void void localJoin_merge_void_void both branches to void
Root merge t t localJoin_merge_tt joins at $t$ (context-lifted)
Root integrate (delta t) localJoin_int_delta unique target void
@@
Root eqW a b, a $\neq$ b localJoin_eqW_ne only diff applies
Root eqW a a, guard ne localJoin_eqW_refl_guard_ne refl blocked; only diff
Ctx merge a b (n-V, n-E) localJoin_ctx_merge_no_void_neq from root ‘merge’ lemma
Ctx eqW a b, a $\neq$ b localJoin_ctx_eqW_ne only diff applies
Ctx eqW a a, guard localJoin_ctx_eqW_refl_guard_ne refl blocked
Ctx eqW a a (norm) localJoin_ctx_eqW_refl_if_norm via normalizeSafe
Table 3: Confluence Lemmas: Verified local-join witnesses for the safe relation (Root and Context).
@@
This summarizes the exact orientation strategy realized in the proofs:
@@
  • The rule rec_zero uses a DM multiset drop on $\kappa^{M}$.
  • The rule rec_succ drops the $\delta$ phase bit (1 $\to$ 0).